a) A wire 40 cm long conducts a current of 2A perpendicular to the direction of a magnetic field B=0.25T. Find the magnetic force over the wire.

b) An alpha particle (+2e) its projected into a magnetic field of 0.14T with a velocity of 3.9x106m/s. Find the magnetic force over the charge at the instant in which the direction of its velocity forms an angle of 55° with respect to the magnetic flux.



1
Expert's answer
2021-11-18T14:59:06-0500

(a) we have the parameters

Length of wire, "l" = 40cm =0.4m

Current, "I" = 2A

Magnetic field, "B" = 0.25T

We also have that, the current is perpendicular to the direction of the magnetic field

Therefore, "\theta""90\u00b0"

We use the formula

"F=BIlsin\u03b8"

"=0.25*2*0.4*sin90\u00b0"

"=0.25*2*0.4*1"(since "Sin90\u00b0=1")

"=0.25*2*0.4"

"=0.25*0.8"

"=0.2""N"



(b) We have the parameters;

Charge, q = +2e

Magnetic field, B =0.14T

Velocity, v = "3.9*10^{-6}" m/s

Angle, "\theta=55°"


Now, we use the magnetic force formula

"F=qvB*sin\theta"

"F=2*3.9*10^{-6}*0.14*sin55\u00b0"

"F=2*3.9*10^{-6}*0.14*0. 8192"

"F=0.229376*3.9*10^{-6}"

"F=8.9457*10^{-7}N"